OpenStudy (anonymous):

The population of a town in 2000 was 430. The population is increasing at a rate of 0.9% every year. What will be the projected population of the town in 2010? Round your answer to the nearest whole number. A. 440 B. 450 C. 460 D. 470

OpenStudy (johnweldon1993):

Population = initial population(1 + r)^n \[P = P_o(1+r)^n\] P_o is the initial population (430) r is the rate at which it is increasing (.009) n is the number of years after 2000 (10) so... \[P = 430(1.90)^{10}\]

OpenStudy (johnweldon1993):

EDIT!*** \[P = 430(1.009)^{10}\]

OpenStudy (anonymous):

i got 470.30

OpenStudy (anonymous):

so D

OpenStudy (johnweldon1993):

That would be correct!